x + y + z =8
x – y – z = -4
-x + y – z = -8

Answer by drj(1380)

x + y + z =8 Equation A
x – y – z = -4 Equation B
-x + y – z = -8 Equation C

Step 1. Add equations A and B where the y and z terms cancel out or eliminated to give:

2x=4 or x=2
Step 2. Add equations B and C where the x and y terms cancel out or eliminated to give:

-2z=-12 or z=6

Step 3. Add equations A and C where the x and z terms cancel out or eliminated to give:

2y=0 or y=0

Step 4. ANSWER: x=2, y=0, and z=6.

Check with original equations which all leads to true statements

x + y + z = 2+0+6 = 8 Equation A
x – y – z = 2-0-6=-4 Equation B
-x + y – z = -2+0-6=-8 Equation C

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J